OpenStudy (anonymous):

Anyone that can help me out? Show me where I am going wrong please? 15x^5+3x+6 ______________ 3x

OpenStudy (anonymous):

any help would be appreciated, i have no idea what i'm doing... well i get ''some'' of it, but I am really confused.

OpenStudy (campbell_st):

well if you split the numerator into 3 parts you get \[\frac{15x^5}{3x} + \frac{3x}{3x} + \frac{6}{3x} = 5x^4 + 1 + \frac{2}{x}\]

OpenStudy (campbell_st):

or if you want everything over the same denominator its \[\frac{5x^4}{1} \times \frac{x}{x} + \frac{x}{x} + \frac{2}{x} = \frac{5x^5 + x + 2}{x}\]

OpenStudy (anonymous):

3x^5-1+4 - x where am I going wrong?

OpenStudy (campbell_st):

ok... so I'll factor the numerator and denominator \[\frac{15x^5 + 3x + 6}{3x} = \frac{3(5x^5 + x +2)}{3(x)} \] 3 is in the numerator(top) and denominator Bottom... 3/3 = 1 so you are left with \[\frac{5x^5 + x + 2}{x}\]